OpenStudy (anonymous):

6.Using complete sentences, explain how to find the zeros of the function f(x) = 2x^3 – 9x + 3. (2 points)

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

What does it mean when it says find the 0's?

OpenStudy (anonymous):

x values that when plugged in, produce a value of zero

OpenStudy (anonymous):

Oh. Okay.

OpenStudy (anonymous):

Thanks.

OpenStudy (lgbasallote):

it means let f(x) = 0 then solve for x

OpenStudy (anonymous):

How would I do that?

OpenStudy (zzr0ck3r):

um thats what they are yes, but I would not put that as your answer. i think they are asking you to show the steps of solving x when y = 0.

OpenStudy (anonymous):

Graph the function and find the x intercepts. OR Factor the equation and solve for f(x) = 0

OpenStudy (lgbasallote):

simple f(x) = 2x^3 + 9x + 2 when f(x) = 0 then 2x^3 + 9x + 2 = 0

OpenStudy (lgbasallote):

factor that out and you get your zeroes

OpenStudy (anonymous):

once you get your values, check http://www.wolframalpha.com/input/?i=2x3+%E2%80%93+9x+%2B+3+%3D0 to see how it works

OpenStudy (anonymous):

Lord victorius use the rational roots theorem and try the possible roots with sythetic division. if you get a remainder of zero then that is one root. according to fundamental theorem of algebra there are 3 zeroes so keep that in mind

OpenStudy (lgbasallote):

do you want a demonstration example?

OpenStudy (zzr0ck3r):

what are you doing in class? The AC method, the quadratic formula, or completing the square?

OpenStudy (anonymous):

AC? .

OpenStudy (anonymous):

Yes, I would like an example.

OpenStudy (lgbasallote):

okay let's say find the zeroes of x^2 + 4x + 3 i just equate it to zero x^2 + 4x + 3 = 0 now i factor it out (x+3)(x+1) = 0 now i equate the factors to 0 x + 3 = 0 so x = -3 x + 1 = 0 so x = -1 does that help?

OpenStudy (anonymous):

Oh! I get it, thanks!

OpenStudy (lgbasallote):

welcome ^_^